OpenStudy (anonymous):

Use a common denominator to determine whether the ratios form a proportion. 6 : 15 and 30 : 80

OpenStudy (anonymous):

What is the greatest common factor of 6 and 15? What is the greatest common factor of 30 and 80?

OpenStudy (anonymous):

3 and 10

OpenStudy (anonymous):

Yes, divide each of the ratio's both sides by the greatest common factor. Divide each side of 6:15 by 3, and divide each side of 30:80 by 10...

OpenStudy (anonymous):

which ratios do you get now? list them, what are they?

OpenStudy (anonymous):

2:5 and 10:8

OpenStudy (anonymous):

the first one is correct, the second on is 3:8.

OpenStudy (anonymous):

\[\frac{2}{5}~~\text{___}~~\frac{3}{8}\] \[\frac{2 \times 8 }{5 \times 8}~~\text{___}~~\frac{3 \times 5 }{8 \times 5}\] \[\frac{16 }{40}~~\text{___}~~\frac{15 }{40}\] fill in the blank with, ≥, ≤ or = .

OpenStudy (anonymous):

Yeah it's >

OpenStudy (anonymous):

I meant > < (sorry it's >)

OpenStudy (anonymous):

So the first ratio is greater:)

OpenStudy (anonymous):

Thank you! :)